The height of a triangle is 8 m less than the base. The area is 10 m squared. Find the height and base.
Answer by dkppathak(439)

The height of a triangle is 8 m less than the base. The area is 10 m squared. Find the height and base.
let base is x m than height will be X-8 m
area of triangle =10m^2 to find x =?
area of triangle =basexheight/2
10=x(x-8)/2
20 =x^2-8X
x^2-8x-20=0
x^2-10x+2x-20=0
(x+2)(x-10)=0
X=10 or x=-2 neglecting -2 so that x=10
base=10m and height will be 2m